Carl bought a backpack with a $82 gift card. The backpack was $x off and the sales tax added $4 to the cost (C). The original price of the backpack was $20.

C = (20 - x) + 4
Remaining money on gift card = 82 - C

Determine how much money Carl has left on the gift card in terms of x.
how do i figure out x?

Answer by stanbon(75887)

Carl bought a backpack with a $82 gift card. The backpack was $x off and the sales tax added $4 to the cost (C). The original price of the backpack was $20.
C = (20 - x) + 4
Remaining money on gift card = 82 - C
Determine how much money Carl has left on the gift card in terms of x.
----
R = 82-C
R = 82-[(20-x)+4]
R = 82-[24-x]
Remainder = 52+x dollars
----------
Cheers,
Stan H.
----------